OpenStudy (anonymous):

Evaluate the expression when p = 9.3 and q = –2.1. p – 3q A. 15.6 B. 3.0 C. –13.2 D. –15.6

OpenStudy (anonymous):

b?

OpenStudy (alexandervonhumboldt2):

9.3-3*(-2.1)=9.3+3*2.1=15.6

OpenStudy (anonymous):

Simplify. 26 + (–3) ∙ (–8) A. –184 B. –2 C. 2 D. 50